A 3.215 kg of unknown substance was transformed from solid to liquid by condensing 10.57 kg of steam under atmospheric conditions.


What is the temperature of steam in degF?

How much heat was gained by the steam in kJ?

What is the heat of fusion of the unknown substance in kJ/g?


1
Expert's answer
2021-11-02T02:23:59-0400

1) 10.57/3.215=3.3•

2) 10.57*3.3=34.75

3) 3.215*3.3=10.6095